For a rectangular garden, the length is 12 feet and the width is 7 feet. Find the length of the diagonal. Round to the nearest tenth of a foot.
Answer by mananth(16949)

The diagonal forms the hypotenuse of the sides.
length = 12
width = 7
..
12^+7^2= diagonal^2
144+49=D^2
193=D^2
=13.9 feet